Leon Inc. has the following capital structure, which it considers to be optimal:

Debt 25%
Preferred Stock 15%
Common Equity 60%

Leon’s expected net income this year is $34,285.72, its established dividend payout ratio is 30%, its federal-plus-state tax rate is 25%, and investors expect future earnings and dividends to grow at a constant rate of 9%. Leon paid a dividend of $3.60 per share last year, and its stock currently sells for $54.00 per share. Leon can obtain new capital in the following ways:

1. New preferred stock with a dividend of $11.00 can be sold to the public at a price of $95.00 per share.

2. Debt can be sold at an interest rate of 12%.

Project Cost at t=0 Rate of Return
A $10,000 17.4%
B $20,000 16%
C $10,000 14.2%
D $20,000 13.2%
E $10,000 12%

Calculate the Retained earnings breakpoint. Assume that Leon does not want to issue any new common stock.
Answer #1

Retained earnings breakpoint = addition to retained earnings/weight of equity in the capital structure

Addition to retained earnings = (1-payout ratio)*net income = (1-30%)*34,285.72 = 24,000.00

Retained earnings break-point = 24,000/60% = 40,000.01